# DisCo: Disentangled Control for Referring Human Dance Generation in Real World

### v1

@@ Appendix A Detailed Discussion on Related Work
Image/Video Editing Conditioned on Text. The most common approach for preserving specific image information is to edit existing images [41, 16, 4, 28, 62] and videos [71, 36, 45, 53] with text, instead of unconditioned generation solely reliant on text descriptions. For example, Prompt-to-Prompt [16] control the spatial layout and geometry of the generated image by modifying the cross-attention maps of the source image. SDEdit [41] corrupts the images by adding noise and then denoises it for editing. DiffEdit [7] first automatically generates the mask highlighting regions to be edited by probing a diffusion model conditioned on different text prompts, then generates edited image guided by the mask. Another line of work requires parameter fine-tuning with user-provided image(s). For example, UniTune [64] tries to fine-tune the large T2I diffusion model on a single image-text pair to encode the semantics into a rare token. The editing image is generated by conditioning on a text prompt containing such rare token. Similarly, Imagic [28] optimizes the text embedding to reconstruct reference image and then interpolate such embedding for image editing.
@@ Appendix B Subject-Specific Finetuning
As mentioned in the main paper, our DisCo can be flexibly integrated with existing efficient fine-tuning techniques for even more fine-grained human dance synthesis. This is particularly beneficial when facing out-of-domain reference images, which appear visually different to the TikTok style images. Figure 6 presents the framework for subject-specific fine-tuning, which is easily adapted from the framework presented in the main text (Figure 2a). Rather than utilizing a set of videos of different human subjects for training, subject-specific fine-tuning aims to leverage limited video frames of a specific human subject (e.g., the video of Elon Mask talking about Tesla Model 3 in Figure 6 or even anime in Figure 7) for better dance synthesis. Compared to the standard fine-tuning, we additionally freeze the pose ControlNet branch and most parts of U-Net to avoid over-fitting to the limited poses in the subject-specific training video, only making the background ControlNet branch and the cross-attention layers in U-Net trainable. We also explored the widely-used LoRA [23] for parameter-efficient fine-tuning and observed similar generation results. As this is not the main focus of this paper, we leave other advanced techniques to future explorations along this direction.
@@ Appendix C Addition Results
DisCo 61.06 0.631 28.78 0.317 4.46E-04 32.56 555.44
DisCo + TikTok HAP 50.68 0.648 28.81 0.309 4.27E-04 30.99 569.45
ControlNet (fg+bg, no SD-VAE) 83.53 0.575 28.37 0.411 5.35E-04 47.69 639.11
ControlNet (fg+bg) 65.14 0.600 28.57 0.355 4.83E-04 33.16 527.74
Attention (fg+bg) 80.50 0.474 28.01 0.485 7.50E-04 80.76 831.10
Ablation on HAP w/ pose (DisCo setting: HAP w/o pose)
TikTok HAP w/ pose 51.84 0.650 28.89 0.307 4.16E-04 29.59 547.14
Ablation on initializing w/ pre-trained ControlNet-Pose [73] (DisCo setting: initialize with U-Net weights)
ControlNet-Pose Init. 62.18 0.633 28.37 0.320 4.46E-04 31.90 591.28
ControlNet-Pose Init.+TikTok HAP 55.81 0.641 28.69 0.316 4.43E-04 30.29 576.67
C.1 Quantitative Results
Figure 8: The effect of different classifier-free-guidance scale.
We show the full ablation results on architecture design in Table 4. In what follows, we focus on discussing results that are not present in the main text. In the first block of the table, we copy over the results from the full instances of DisCo, with or without HAP on TikTok Dance Dataset for reference.
As mentioned in L176-178 of the main text, we propose to use the pre-trained VQ-VAE from SD, instead of four randomly initialized convolution layers in the original ControlNet for encoding the background reference image. In the second block of the table, we ablate this design by comparing two models, (1) “ControlNet (fg+bg)”, inserting the control condition of reference image (containing both foreground and background) via a single ControlNet, with VQ-VAE encoding and (2) “ControlNet (fg+bg, no SD-VAE)”, inserting the control condition of reference image via a single ControlNet with four randomly initialized convolution layers as the condition encoder. We note that the pre-trained VQ-VAE can produce a more descriptive dense representation of the reference image, contributing to better synthesis results (FID $65.14$ v.s $83.59$).
In the third block of the table, we investigate whether adding pose condition into human attribute pretraining is beneficial to the downstream performance. We observe that integrating pose into HAP leads in similar results, but requires additional annotation efforts on pose estimation.
Last but not least, we examine on the initialization of the pose ControlNet branch. Specifically, we try to initialize from the pre-trained ControlNet-Pose checkpoint [73] during fine-tuning. The results are shown in the last block of Table 4. Without HAP, the performance is comparable to DisCo, but it gets much worse than DisCo when both are pre-trained with HAP. This is because that the ControlNet-Pose is pre-trained with text condition and can not fully accommodate referring human dance generation with the reference image condition. After HAP, such gap is further enlarged, leading to even worse results. In Figure 8, we show the effect of varying the classifier-free-guidance scale. We can find that scale of $1.5$ gives the best quantitative results for both image-wise and video-wise fidelity.
@@ C.1 Quantitative Results
Figure 9: More visualizations for different scenarios of human image editing.
C.2 Qualitative Results
Figure 9 presents additional results from DisCo for different human image editing scenarios. Similar to Figure 3 in the main text, DisCo can handle novel human pose synthesis very well even with change in viewpoints and rotation in the human skeleton. More results for video generation is shown in Figure 11.

### v3

@@ Appendix A Detailed Discussion on Related Work
Image/Video Editing Conditioned on Text. The most common approach for preserving specific image information is to edit existing images [38, 14, 26, 59] and videos [68, 35, 42, 50] with text, instead of unconditioned generation solely reliant on text descriptions. For example, Prompt-to-Prompt [14] control the spatial layout and geometry of the generated image by modifying the cross-attention maps of the source image. SDEdit [38] corrupts the images by adding noise and then denoises it for editing. DiffEdit [6] first automatically generates the mask highlighting regions to be edited by probing a diffusion model conditioned on different text prompts, then generates edited image guided by the mask. Another line of work requires parameter fine-tuning with user-provided image(s). For example, UniTune [61] tries to fine-tune the large T2I diffusion model on a single image-text pair to encode the semantics into a rare token. The editing image is generated by conditioning on a text prompt containing such rare token. Similarly, Imagic [26] optimizes the text embedding to reconstruct reference image and then interpolate such embedding for image editing.
@@ Appendix B Subject-Specific Finetuning
DisCo 61.06 0.631 15.38 0.317 4.46E-04 73.29 366.39
DisCo + TikTok HAP 50.68 0.648 15.62 0.309 4.27E-04 69.68 353.35
ControlNet (fg+bg, no SD-VAE) 83.53 0.575 14.51 0.411 5.35E-04 89.13 551.62
ControlNet (fg+bg) 65.14 0.600 15.15 0.355 4.83E-04 74.19 427.49
Attention (fg+bg) 80.50 0.474 12.38 0.485 7.50E-04 80.49 551.51
Ablation on HAP w/ pose (DisCo setting: HAP w/o pose)
TikTok HAP w/ pose 51.84 0.650 15.73 0.307 4.16E-04 68.55 346.10
Ablation on initializing w/ pre-trained ControlNet-Pose (DisCo setting: initialize with U-Net weights)
ControlNet-Pose Init. 62.18 0.633 15.42 0.320 4.46E-04 72.98 389.47
ControlNet-Pose Init.+TikTok HAP 55.81 0.641 15.37 0.316 4.43E-04 78.13 363.38
As mentioned in the main paper, our DisCo can be flexibly integrated with existing efficient fine-tuning techniques for even more fine-grained human dance synthesis. This is particularly beneficial when facing out-of-domain reference images, which appear visually different to the TikTok style images. Figure 8 presents the framework for subject-specific fine-tuning, which is easily adapted from the framework presented in the main text (Figure 0(a)). Rather than utilizing a set of videos of different human subjects for training, subject-specific fine-tuning aims to leverage limited video frames of a specific human subject (e.g., the video of Elon Mask talking about Tesla Model 3 in Figure 8 or even anime in Figure 9) for better dance synthesis. Compared to the standard fine-tuning, we additionally freeze the pose ControlNet branch and most parts of U-Net to avoid over-fitting to the limited poses in the subject-specific training video, only making the background ControlNet branch and the cross-attention layers in U-Net trainable. We also explored the widely-used LoRA [21] for parameter-efficient fine-tuning and observed similar generation results. As this is not the main focus of this paper, we leave other advanced techniques to future explorations along this direction.
Figure 10: The effect of different classifier-free-guidance scale.
@@ Appendix C Addition Results
C.1 Adaptable to Other Conditions
DisCo can be easily extended to face landmarks and hand gestures to provide more fine-grained control.
C.2 More Quantitative Results
@@
Figure 11: More visualizations for different image ratios (i.e., vertical image) and various human views (e.g., from close-view to full-body) of human image editing.
Figure 13: The interface of our user study. We provide detailed instructions, rating examples, and rating instances for users.
Full Ablation Results. We show the full ablation results on architecture design in Table 6. In what follows, we focus on discussing results that are not present in the main text. In the first block of the table, we copy over the results from the full instances of DisCo, with or without HAP on TikTok Dance Dataset for reference.
As mentioned in the main text, we propose to use the pre-trained VQ-VAE from SD, instead of four randomly initialized convolution layers in the original ControlNet for encoding the background reference image. In the second block of the table, we ablate this design by comparing two models, (1) “ControlNet (fg+bg)”, inserting the control condition of reference image (containing both foreground and background) via a single ControlNet, with VQ-VAE encoding and (2) “ControlNet (fg+bg, no SD-VAE)”, inserting the control condition of reference image via a single ControlNet with four randomly initialized convolution layers as the condition encoder. We note that the pre-trained VQ-VAE can produce a more descriptive dense representation of the reference image, contributing to better synthesis results (FID $65.14$ v.s $83.59$). In the third block of the table, we investigate whether adding pose condition into human attribute pretraining is beneficial to the downstream performance. We observe that integrating pose into HAP leads in similar results, but requires additional annotation efforts on pose estimation.
Last but not least, we examine on the initialization of the pose ControlNet branch. Specifically, we try to initialize from the pre-trained ControlNet-Pose checkpoint [71] during fine-tuning. The results are shown in the last block of Table 6. Without HAP, the performance is comparable to DisCo, but it gets much worse than DisCo when both are pre-trained with HAP. This is because that the ControlNet-Pose is pre-trained with text condition and can not fully accommodate referring human dance generation with the reference image condition. After HAP, such gap is further enlarged, leading to even worse results.
Effect of CFG Scale. In Figure 10, we show the effect of varying the classifier-free-guidance scale. We can find that scale of $1.5$ gives the best quantitative results for both image-wise and video-wise fidelity.
Figure 14: Full results of User Study: (a) human ID alignment; (b) pose alignment; and (c) background alignment score distribution on both synthesis images (above part) and videos (below part) of DreamPose and DisCo.
Pre-trained Motion Prior. In addition to utilizing the appearance pre-training, we also investigate the potential of the motion prior learned from the large-scale general video. We inject motion prior from the AnimateDiff (AD) [13] checkpoint pre-trained on the large-scale video and then fine-tune on TikTok data.
Method FID-VID $\downarrow$ FVD $\downarrow$
w/o. AD 20.75 257.90
w/. AD 18.96 241.55
Results in left table show superior temporal smoothness, making it worthwhile to further explore motion pre-training via video data together with attribute pre-training via image corpus.
C.3 Qualitative Results
DisCo can be easily adapted to different image size. For example, we show more qualitative results of human image editing in Figure 11 with image size of $256\times 384$ to include more human body. Please note that most videos of the TikTok dataset are relatively close to the camera. However, we can see that DisCo can handle both partial and full human body synthesis even with large changes in viewpoints and rotations in the human skeleton. More results for video generation are shown in Figure 15.
@@ C.3 Qualitative Results
C.4 User Study
As indicated in the main paper, we conducted a user study involving $50$ distinct individuals to compare our DisCo with its key counterpart, DreamPose [25] to evaluate the quality of the generated images and videos. The study contains $80$ rating questions for $20$ synthesis ($10$ images and $10$ videos) in total. There are $5$ images and $5$ videos for both DisCo and DreamPose. In addition to the coarse-grained overall quality score adopted in DreamPose, we further divide the rating evaluation into three fine-grained aspects: (i) human ID alignment; (ii) pose alignment; and (iii) background alignment. For each aspect, the users are required to rate on a scale of 00 to $5$, where 00 corresponds to an ideal corresponding (e.g., the ground-truth image/video) to the reference input image and 00 for non-match at all. This results in $4000$ responses in total. Please check Figure 13 for the user study interface and Figure 14 for the full results.
We can observe that, for both human ID and background, DisCo (green histogram) achieves clearly higher scores compared to DreamPose (red one). This indicates that the naive combination of different conditions (e.g., DreamPose) suffers from inconsistent human attribute and unstable background, while DisCo can reconstruct the fine-grained human attributes and main steady backgrounds thanks to the proposed human attribute pre-training and well-designed disentangled control. As a much easier condition, pose control can be generally maintained for both DreamPose and DisCo. However, DisCo still demonstrates obvious superiority compared to DreamPose.
C.5 Ethical Concern
Despite the broad applicability of DisCo, it is crucial to consider the risks of misuse in creating deceptive media, address potential biases in training data, and ensure respect for intellectual property.